Old School Publishing Inc. began printing operations on January 1. Jobs 301 and 302 were completed...

  1. Old School Publishing Inc. began printing operations on January 1. Jobs 301 and 302 were completed during the month, and all costs applicable to them were recorded on the related cost sheets. Jobs 303 and 304 are still in process at the end of the month, and all applicable costs except factory overhead have been recorded on the related cost sheets. In addition to the materials and labor charged directly to the jobs, $2,910 of indirect materials and $35,210 of indirect labor were used during the month. The cost sheets for the four jobs entering production during the month are as follows, in summary form:

    Job 301 Job 302
    Direct materials $30,110 Direct materials $14,080
    Direct labor 11,500 Direct labor 6,000
    Factory overhead 6,325 Factory overhead 3,300
    Total $47,935 Total $23,380
    Job 303 Job 304
    Direct materials $42,300 Direct materials $8,670
    Direct labor 12,800 Direct labor 1,700
    Factory overhead Factory overhead

    Journalize the summary entry to record each of the following operations for January (one entry for each operation):

    a. Direct and indirect materials used. For a compound transaction, if an amount box does not require an entry, leave it blank.

    Work in Process
    • Cash
    • Cost of Goods Sold
    • Factory Expenses
    • Finished Goods
    • Materials
    • Work in Process
    Factory Overhead
    • Accounts Payable
    • Cash
    • Cost of Goods Sold
    • Factory Overhead
    • Finished Goods
    • Materials
    Materials
    • Accounts Payable
    • Cash
    • Cost of Goods Sold
    • Factory Overhead
    • Materials
    • Work in Process

    Feedback

    b. Direct and indirect labor used. For a compound transaction, if an amount box does not require an entry, leave it blank.

    • Accounts Payable
    • Finished Goods
    • Materials
    • Wages Payable
    • Work in Process
    • Accounts Payable
    • Cash
    • Factory Overhead
    • Materials
    • Wages Payable
    • Accounts Payable
    • Factory Overhead
    • Materials
    • Wages Payable
    • Work in Process

    Feedback

    c. Factory overhead applied to all four jobs (a single overhead rate is used based on direct labor cost).

    • Accounts Payable
    • Cash
    • Factory Overhead
    • Materials
    • Work in Process
    • Accounts Payable
    • Cash
    • Factory Overhead
    • Finished Goods
    • Work in Process

    Feedback

    d. Completion of Jobs 301 and 302.

    • Cash
    • Factory Overhead
    • Finished Goods
    • Materials
    • Work in Process
Answer #1
a
Work in Process 95160 =30110+14080+42300+8670
Factory Overhead 2910
        Materials 98070
b
Work in Process 32000 =11500+6000+12800+1700
Factory Overhead 35210
        Wages Payable 67210
c
Work in Process 17600 =32000*55%
       Factory Overhead 17600
d
Finished Goods 71315 =47935+23380
       Work in Process 71315